Evaluate
(-4x)small 2 and -4x small2 when x = 4
I did not know how to put the small 2 next to the number sorry

Found 2 solutions by rapaljer, stanbon:
Answer by rapaljer(4671)

Is the small 2 written as a superscript (an exponent?) or a subscript??? My guess is that these are exponents!

R^2

Answer by stanbon(75887)

(-4x)small 2 and -4x small2 when x = 4
--------------------
(-4x)^2 when x = 4 is (-4*4)^2 = (-16)^2 = 256
-------------------
-4x^2 when x = 4 is -4*4^2 = -4*16 = -256
=======================
Cheers,
Stan H.